# Peaceful Anarcho-Accelerationism: Decentralized Full Automation for a Society of Universal Care

### v5

@@ Abstract
Foundational results in machine learning establish that all human labor is in principle automatable. This paper introduces peaceful anarcho-accelerationism, a rigorously defined sociotechnical framework grounded in the 200-year anarchist tradition from Godwin through Kropotkin to Bookchin, and in the methodological categories of Eltzbacher, Nettlau, and Corrêa, for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. We state five formal hypotheses and six research objectives, present a formal definition through analytical categories of interdependent spheres, and propose the Liberation Stack as a layered technical architecture with explicit preconditions and gate conditions for each layer. We introduce Universal Desired Resources as a post-monetary design principle that eliminates the material basis of intersectional oppression, and develop a framework for progressive state dissolution through incremental, reversible commons-building compatible with existing democratic institutions. A nonviolent social mobilization strategy maps concrete peaceful methods to each stage of transition. We show that accelerationism and degrowth share anarchist pacifism as substrate and differ only along a Pareto-optimal technological frontier. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, guifi.net, the Fediverse, and contemporary commons initiatives confirms that commons-based systems already operate at scale. We conclude with a phased roadmap specifying explicit assumptions, hard constraints, gate conditions between phases, and detailed limitations.
@@ 1 The Convergence Window and the Automation Imperative
This is not a resource problem. It is an architecture problem.
Before proceeding, we state the formal hypotheses that this paper investigates and the research objectives it pursues, so that the reader can evaluate the argument against explicit claims rather than vague aspirations.
The first hypothesis (H1) is that all human labor, both cognitive and physical, is in principle automatable. This claim rests on the universal approximation theorem (Cybenko, 1989; Hornik, 1991) and the convergence results of deep reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2015; Watkins and Dayan, 1992), which together establish that no mathematical barrier prevents the automation of any formalizable task or any learnable sensorimotor policy. We emphasize that this is an in-principle claim: the existence of approximating networks and convergent policies does not guarantee tractable training, nor does it address tasks whose specification may be inherently incomplete. The second hypothesis (H2) is that decentralized commons governance of automation produces better outcomes than centralized corporate ownership, measured along three dimensions: inequality (Gini coefficient and wealth distribution), human well-being (life satisfaction, health outcomes, educational attainment, and Maslow-level satisfaction), and ecological footprint (carbon emissions, resource throughput, and biodiversity impact). This hypothesis is testable through comparative analysis of existing commons-based systems and through pilot implementations of the Liberation Stack. The third hypothesis (H3) is that nonviolent social mobilization can achieve the commons transition without state seizure, drawing on the empirical findings of Chenoweth and Stephan (2011) that nonviolent campaigns succeed at twice the rate of violent ones, and on the strategic framework of Sharp (1973) that identifies 198 methods of nonviolent action applicable to the commons transition. The fourth hypothesis (H4) is that the apparent opposition between degrowth and accelerationism is a false binary: both positions share anarchist pacifism as their ethical substrate and occupy different points along a Pareto-optimal frontier trading off ecological risk against human suffering during the transition. A mature movement would support both strategies simultaneously, and technological progress shifts the entire frontier toward the origin. The fifth hypothesis (H5) is that each layer of the Liberation Stack has identifiable preconditions for successful implementation, and that the layers form a partial ordering of dependencies: energy is the precondition for all material layers, communication is the precondition for knowledge and governance, and no layer can be sustainably deployed without the layers below it in the dependency graph.
@@
The urgency of this framework stems from a technological reality that is no longer speculative: full automation of human labor is theoretically possible in principle and empirically underway. The universal approximation theorem, established independently by Cybenko (Cybenko, 1989) and Hornik et al. (Hornik, 1991), proves that a feedforward neural network with a single hidden layer containing a finite number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary precision. Formally, for any continuous function $f:K\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$ and any $\epsilon>0$, there exists a neural network $\hat{f}$ such that $\|f(x)-\hat{f}(x)\|<\epsilon$ for all $x\in K$. We note important caveats: the UAT is an existence result that guarantees an approximating network exists but provides no constructive procedure for finding it, no bound on the network size required, and no guarantee that gradient-based training will converge to the approximator in tractable time. Our claim is therefore that full automation is possible in principle: the theoretical barrier has been removed, not that it is trivially achievable for all tasks.
@@
If both cognitive and physical labor are fully automatable, then the wage-labor system that has organized human societies for two centuries faces dissolution. Humanity confronts a fork, illustrated in Figure˜2. On one path, robots and AI systems remain owned by corporations; employment disappears but ownership stays concentrated; the majority of humans become economically superfluous, dependent on corporate benevolence or state transfers for survival. This is neo-feudalism with a technological veneer (Zuboff, 2019). On the other path, robots and AI systems are governed as commons, owned by communities, operated transparently, and maintained collectively. Automated production provides universal access to goods and services. Humans, freed from instrumental labor, dedicate their time to what machines cannot and should not replace: caring for each other, raising children, building community, creating art, pursuing knowledge, loving.
@@
Path B is a care-centered society: a society where the fundamental organizing principle is not production for profit but care for persons. The implications of full automation extend further than this binary choice. The first consequence is the obsolescence of money. Money exists as a coordination mechanism for allocating scarce goods produced by human labor. When all production is automated and the marginal cost of most goods approaches zero, the function of money dissolves. We propose the concept of Universal Desired Resources (UDR): in a fully automated commons, every person has direct access to the goods and services they desire, requested from the robotic productive commons, subject only to ecological sustainability constraints. The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The third consequence is an ethical clarification: it is not only permissible but morally required to delegate all instrumental labor to machines, on the basis of a clear ontological distinction between computational intelligence and phenomenal consciousness (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022). A robot that assembles furniture does not feel boredom. An AI that writes code does not feel alienation. Delegating labor to entities that cannot suffer, in order to liberate entities that can, is care applied at civilizational scale. The fourth consequence is that morality itself becomes computationally tractable: human moral decision-making can be modeled as a Markov Decision Process where the reward function encodes humanist values (Garrido-Merchán and Lumbreras-Sancho, 2023), though this formalization requires strong assumptions about value alignment and moral stationarity.
@@
These movements failed not because their organizational principles were flawed, but because they lacked three conditions that are now, for the first time in history, within reach. First, they had no technological infrastructure for decentralized coordination at scale: no internet, no mesh networks, no federated protocols, no real-time communication across territories. The CNT’s extraordinary organizational achievement of reaching 1.5 million members relied on physical meetings, printed newspapers, and postal correspondence, all vulnerable to disruption by military force. Today, federated communication protocols (ActivityPub, Matrix), community mesh networks (guifi.net), and open-source governance tools (Decidim, Loomio) provide the infrastructure for coordination that the historical movements could only dream of. Second, they had no energy independence. Every historical anarchist territory depended on external energy sources controlled by hostile states. Community-owned solar microgrids, enhanced geothermal systems, and iron-air batteries now make energy sovereignty technically and economically feasible at the community scale. Third, they had no productive automation. The collectives of 1936 still required human labor for every productive task, making them vulnerable to the economic disruptions of war. Autonomous robotic systems for agriculture, manufacturing, and logistics now make economic self-sufficiency feasible without requiring the full participation of every community member in manual labor.
The remainder of this paper provides the formal definition (Section˜2), the political bridges (Section˜3), the technical architecture and ecological framework (Section˜4), the nonviolent mobilization strategy (Section˜5), the empirical evidence (Section˜6), and the roadmap (Section˜7) for building this society.
@@ 2 Theoretical Foundations: A Formal Definition of Peaceful Anarcho-Accelerationism
Anarcho-accelerationism does not emerge from a vacuum. It is grounded in a 200-year intellectual tradition of anarchist political philosophy, enriched by the systematic historiography of scholars who have mapped this tradition with increasing rigor. Our methodology draws on six major historians of anarchism and updates their analytical categories for the technological conditions of the 21st century, following the principle of totality and interdependence articulated by Corrêa (Corrêa, 2015).
Paul Eltzbacher (Eltzbacher, 1908) produced the first systematic comparative analysis of anarchist thought, examining seven thinkers, namely Godwin, Proudhon, Stirner, Bakunin, Kropotkin, Tucker, and Tolstoy, through a rigorous analytical framework that identified their positions on law, the state, and property. Eltzbacher’s method of systematic comparison across shared analytical categories remains foundational to our approach. Max Nettlau (Nettlau, 1996), known as the Herodotus of Anarchy, produced the most comprehensive historiographic corpus in anarchist studies, documenting movements across continents and centuries in his multi-volume Geschichte der Anarchie. Nettlau’s contribution is methodological: he demonstrated that anarchism is not a singular doctrine but a family of movements unified by shared opposition to domination and commitment to voluntary association. George Woodcock (Woodcock, 1962) provided the standard English-language history, periodizing anarchist thought from Godwin’s Enquiry Concerning Political Justice (Godwin, 1793) through the classical period to the Spanish Revolution and beyond. James Joll (Joll, 1964) situated anarchism within the broader context of European political movements. Daniel Guérin (Guérin, 1970) synthesized anarchist and Marxist insights, emphasizing anarchism as a practical movement oriented toward workers’ self-management. Peter Marshall (Marshall, 1992) produced the most comprehensive single-volume history, tracing libertarian ideas from ancient Taoism through the Enlightenment to contemporary movements.
@@
Following the methodology of Corrêa (Corrêa, 2015), who insists on the principle of totality and interdependence in analyzing anarchist movements, we define peaceful anarcho-accelerationism through seven analytical categories that are not intelligible in isolation but condition and are conditioned by each other in a dialectical whole. The seven categories, compared with their classical formulations, are presented in Table˜4.
@@
Corrêa further identifies social vectors, the specific social forces that carry the potential for transformation, and critiques, the specific structures of domination against which the movement is directed. In our framework, the social vectors are the precariat (the growing class of workers in precarious, platform-mediated employment), the open-source development community (which already practices commons-based peer production at planetary scale), the cooperative movement (from Mondragon to platform cooperatives), the agroecological movement, the community energy movement, and the Fediverse community (which builds and maintains decentralized communication infrastructure). The critiques are directed against techno-feudalism (the concentration of digital productive capacity in a handful of platform corporations), surveillance capitalism (the extraction and monetization of behavioral data), algorithmic management (the replacement of human supervisors with opaque automated systems), fossil-fuel centralization (the structural dependence on concentrated energy sources that enables authoritarian control), and intellectual property enclosure (the artificial scarcification of knowledge, seeds, and code through patent and copyright regimes).
@@
Drawing on Kropotkin’s naturalistic ethics (Kropotkin, 1924), Tolstoy’s principled nonviolence (Tolstoy, 1894), Gandhi’s satyagraha (Gandhi, 1961), and the care ethics tradition (Tronto, 1993), we articulate ten ethical principles that govern the peaceful anarcho-accelerationist project. These principles are not abstract ideals but design constraints for the Liberation Stack, and every technology, protocol, and institution proposed in this paper can be evaluated against them. The first is non-domination: no person, institution, or algorithm shall exercise coercive power over another, and technology must distribute power, never concentrate it. The second is universal care: every system, institution, and technology shall be evaluated by its capacity to sustain life, relationships, and communities, since care is the organizing principle, not profit. The third is nonviolence: social transformation shall be achieved through peaceful means, including technological commons-building, education, mutual aid, prefigurative politics, and democratic persuasion, following Gandhi (Gandhi, 1961) and Sharp (Sharp, 1973). The fourth is commons governance: essential resources such as energy, food, knowledge, communication, code, and productive infrastructure shall be governed as commons according to Ostrom’s principles (Ostrom, 1990), not enclosed as property. The fifth is ecological integrity: all productive activity shall operate within planetary boundaries (Raworth, 2017), since sustainability is a first-order constraint, not an externality. The sixth is radical transparency: algorithms that affect human lives shall be auditable, data collection shall be consensual, and governance shall be open, because opacity is a tool of domination. The seventh is solidarity across difference: following intersectional analysis (Crenshaw, 1989; Collins, 2019), liberation must be universal, since no one is free until everyone is free. The eighth is intellectual commitment: following Chomsky’s principle of the responsibility of intellectuals (Chomsky, 1967), those with knowledge and privilege have a heightened obligation to speak truth to power and to build alternatives. The ninth is freedom of conscience: every person’s spiritual, philosophical, and existential commitments shall be respected, since the framework serves human flourishing in all its dimensions, including the spiritual (Lumbreras and Garrido-Merchán, 2024). The tenth is prefigurative consistency: the means must embody the ends, systems designed for democracy must be governed democratically, and tools for liberation must not reproduce domination (Ward, 1973).
@@
Anarcho-accelerationism rests on four intellectual pillars that bridge the historical and the contemporary. Kropotkin’s Mutual Aid (Kropotkin, 1902) challenged the Social Darwinist misreading of natural selection by documenting extensive cooperation across species. Contemporary evolutionary biology has vindicated this view: multilevel selection theory (Wilson, 2007), reciprocal altruism (Trivers, 1971), and cultural group selection (Henrich, 2016) all demonstrate that cooperation is a primary driver of complex social organization. In computer science, this principle is realized in commons-based peer production (Benkler, 2006). Ostrom’s institutional analysis (Ostrom, 1990) demolished Hardin’s tragedy of the commons by documenting thousands of cases where communities successfully manage shared resources without either privatization or state control. Her eight design principles provide a rigorous institutional framework that maps directly onto distributed system design (De Moor, 2010). Following Stiegler (Stiegler, 2010), we adopt the concept of the pharmakon: every technology is simultaneously remedy and poison. This is not a reason for technological pessimism; it is a reason for technological governance (Feenberg, 2002). Drawing from the anarchist tradition of prefigurative politics (Graeber, 2013), we propose that the systems we build should embody the values they aim to promote.

### v6

@@ Abstract
Foundational results in machine learning establish that all human labor may in principle be automatable. Without deliberate intervention, this trajectory risks concentrating productive capacity in a handful of corporations, resulting in techno-feudalism: mass economic redundancy, surveillance-based control, and dependence on corporate benevolence for survival. To avert this outcome, this paper introduces peaceful anarcho-accelerationism, a rigorously defined sociotechnical framework grounded in the 200-year anarchist tradition from Godwin through Kropotkin to Bookchin, and in the methodological categories of Eltzbacher, Nettlau, and Corrêa, for ensuring that full automation is decentralized, commons-governed, and oriented toward universal care. Specifically, I state five formal hypotheses and six research objectives, present a formal definition through analytical categories of interdependent spheres, and propose the Liberation Stack as a layered technical architecture with explicit preconditions and gate conditions for each layer. Furthermore, I introduce Universal Desired Resources as a post-monetary design principle that eliminates the material basis of intersectional oppression, and develop a framework for progressive state dissolution through incremental, reversible commons-building compatible with existing democratic institutions. Additionally, a nonviolent social mobilization strategy maps concrete peaceful methods to each stage of transition. Moreover, I show that accelerationism and degrowth share anarchist pacifism as substrate and differ only along a Pareto-optimal technological frontier. Empirical evidence from Linux, Wikipedia, Mondragon, Rojava, guifi.net, the Fediverse, and contemporary commons initiatives confirms that commons-based systems already operate at scale. Finally, I conclude with a phased roadmap specifying explicit assumptions, hard constraints, gate conditions between phases, and detailed limitations.
@@ 1 The Convergence Window and the Automation Imperative
This disparity suggests that the central challenge is not one of resource scarcity but of sociotechnical architecture: the systems governing production, distribution, and access are structured to concentrate rather than distribute.
Before proceeding, I state the formal hypotheses that this paper investigates and the research objectives it pursues, so that the reader can evaluate the argument against explicit claims rather than vague aspirations.
The first hypothesis (H1) is that all human labor, both cognitive and physical, is in principle automatable. This claim rests on the universal approximation theorem (Cybenko, 1989; Hornik, 1991) and the convergence results of deep reinforcement learning (Sutton and Barto, 2018; Mnih et al., 2015; Watkins and Dayan, 1992), which together establish that no mathematical barrier prevents the automation of any formalizable task or any learnable sensorimotor policy. I emphasize that this is an in-principle claim: the existence of approximating networks and convergent policies does not guarantee tractable training, nor does it address tasks whose specification may be inherently incomplete. The second hypothesis (H2) is that decentralized commons governance of automation produces better outcomes than centralized corporate ownership, measured along three dimensions: inequality (Gini coefficient and wealth distribution), human well-being (life satisfaction, health outcomes, educational attainment, and Maslow-level satisfaction), and ecological footprint (carbon emissions, resource throughput, and biodiversity impact). This hypothesis is testable through comparative analysis of existing commons-based systems and through pilot implementations of the Liberation Stack. The third hypothesis (H3) is that nonviolent social mobilization can achieve the commons transition without state seizure, drawing on the empirical findings of Chenoweth and Stephan (2011) that nonviolent campaigns succeed at twice the rate of violent ones, and on the strategic framework of Sharp (1973) that identifies 198 methods of nonviolent action applicable to the commons transition. The fourth hypothesis (H4) is that the apparent opposition between degrowth and accelerationism is a false binary: both positions share anarchist pacifism as their ethical substrate and occupy different points along a Pareto-optimal frontier trading off ecological risk against human suffering during the transition. A mature movement would support both strategies simultaneously, and technological progress shifts the entire frontier toward the origin. The fifth hypothesis (H5) is that each layer of the Liberation Stack has identifiable preconditions for successful implementation, and that the layers form a partial ordering of dependencies: energy is the precondition for all material layers, communication is the precondition for knowledge and governance, and no layer can be sustainably deployed without the layers below it in the dependency graph.
@@
The urgency of this framework stems from a technological reality that is no longer speculative: full automation of human labor is theoretically possible in principle and empirically underway. The universal approximation theorem, established independently by Cybenko (Cybenko, 1989) and Hornik et al. (Hornik, 1991), proves that a feedforward neural network with a single hidden layer containing a finite number of neurons can approximate any continuous function on compact subsets of $\mathbb{R}^{n}$ to arbitrary precision. Formally, for any continuous function $f:K\subset\mathbb{R}^{n}\to\mathbb{R}^{m}$ and any $\epsilon>0$, there exists a neural network $\hat{f}$ such that $\|f(x)-\hat{f}(x)\|<\epsilon$ for all $x\in K$. I note important caveats: the UAT is an existence result that guarantees an approximating network exists but provides no constructive procedure for finding it, no bound on the network size required, and no guarantee that gradient-based training will converge to the approximator in tractable time. My claim is therefore that full automation is possible in principle: the theoretical barrier has been removed, not that it is trivially achievable for all tasks.
@@
If both cognitive and physical labor are fully automatable, then the wage-labor system that has organized human societies for two centuries faces dissolution. Humanity confronts a fork, illustrated in Figure˜2. On one path, robots and AI systems remain owned by corporations; employment disappears but ownership stays concentrated; the majority of humans become economically superfluous, dependent on corporate benevolence or state transfers for survival. This is neo-feudalism with a technological veneer (Zuboff, 2019). On the other path, robots and AI systems are governed as commons, owned by communities, operated transparently, and maintained collectively. Automated production provides universal access to goods and services. Humans, freed from instrumental labor, dedicate their time to what machines cannot and should not replace: caring for each other, raising children, building community, creating art, and pursuing knowledge.
@@
Path B is a care-centered society: a society where the fundamental organizing principle is not production for profit but care for persons. The implications of full automation extend further than this binary choice. The first consequence is the obsolescence of money. Money exists as a coordination mechanism for allocating scarce goods produced by human labor. When all production is automated and the marginal cost of most goods approaches zero, the function of money dissolves. I propose the concept of Universal Desired Resources (UDR): in a fully automated commons, every person has direct access to the goods and services they desire, requested from the robotic productive commons, subject only to ecological sustainability constraints. The second consequence is the convergence of manufacturing, biology, and intelligence toward effectively unlimited productive capacity. The third consequence is an ethical clarification: it is not only permissible but morally required to delegate all instrumental labor to machines, on the basis of a clear ontological distinction between computational intelligence and phenomenal consciousness (Garrido-Merchán and Lumbreras, 2022; Garrido-Merchán, 2024; Garrido-Merchán and Blanco, 2022). A robot that assembles furniture does not feel boredom. An AI that writes code does not feel alienation. Delegating labor to entities that cannot suffer, in order to liberate entities that can, constitutes a systematic application of care ethics to the organization of production. The fourth consequence is that morality itself becomes computationally tractable: human moral decision-making can be modeled as a Markov Decision Process where the reward function encodes humanist values (Garrido-Merchán and Lumbreras-Sancho, 2023), though this formalization requires strong assumptions about value alignment and moral stationarity.
@@
These movements failed not because their organizational principles were flawed, but because they lacked three conditions that are now, for the first time in history, within reach. First, they had no technological infrastructure for decentralized coordination at scale: no internet, no mesh networks, no federated protocols, no real-time communication across territories. The CNT’s extraordinary organizational achievement of reaching 1.5 million members relied on physical meetings, printed newspapers, and postal correspondence, all vulnerable to disruption by military force. Today, federated communication protocols (ActivityPub, Matrix), community mesh networks (guifi.net), and open-source governance tools (Decidim, Loomio) provide the infrastructure for coordination that the historical movements did not have access to. Second, they had no energy independence. Every historical anarchist territory depended on external energy sources controlled by hostile states. Community-owned solar microgrids, enhanced geothermal systems, and iron-air batteries now make energy sovereignty technically and economically feasible at the community scale. Third, they had no productive automation. The collectives of 1936 still required human labor for every productive task, making them vulnerable to the economic disruptions of war. Autonomous robotic systems for agriculture, manufacturing, and logistics now make economic self-sufficiency feasible without requiring the full participation of every community member in manual labor.
The remainder of this paper provides the formal definition (Section˜2), the political bridges (Section˜3), the technical architecture and ecological framework (Section˜4), the nonviolent mobilization strategy (Section˜5), the empirical evidence (Section˜6), and the roadmap (Section˜7) for the framework proposed in this paper.
@@ 2 Theoretical Foundations: A Formal Definition of Peaceful Anarcho-Accelerationism
Anarcho-accelerationism does not emerge from a vacuum. It is grounded in a 200-year intellectual tradition of anarchist political philosophy, enriched by the systematic historiography of scholars who have mapped this tradition with increasing rigor. My methodology draws on six major historians of anarchism and updates their analytical categories for the technological conditions of the 21st century, following the principle of totality and interdependence articulated by Corrêa (Corrêa, 2015).
Paul Eltzbacher (Eltzbacher, 1908) produced the first systematic comparative analysis of anarchist thought, examining seven thinkers, namely Godwin, Proudhon, Stirner, Bakunin, Kropotkin, Tucker, and Tolstoy, through a rigorous analytical framework that identified their positions on law, the state, and property. Eltzbacher’s method of systematic comparison across shared analytical categories remains foundational to my approach. Max Nettlau (Nettlau, 1996), known as the Herodotus of Anarchy, produced the most comprehensive historiographic corpus in anarchist studies, documenting movements across continents and centuries in his multi-volume Geschichte der Anarchie. Nettlau’s contribution is methodological: he demonstrated that anarchism is not a singular doctrine but a family of movements unified by shared opposition to domination and commitment to voluntary association. George Woodcock (Woodcock, 1962) provided the standard English-language history, periodizing anarchist thought from Godwin’s Enquiry Concerning Political Justice (Godwin, 1793) through the classical period to the Spanish Revolution and beyond. James Joll (Joll, 1964) situated anarchism within the broader context of European political movements. Daniel Guérin (Guérin, 1970) synthesized anarchist and Marxist insights, emphasizing anarchism as a practical movement oriented toward workers’ self-management. Peter Marshall (Marshall, 1992) produced the most comprehensive single-volume history, tracing libertarian ideas from ancient Taoism through the Enlightenment to contemporary movements.
@@
Following the methodology of Corrêa (Corrêa, 2015), who insists on the principle of totality and interdependence in analyzing anarchist movements, I define peaceful anarcho-accelerationism through seven analytical categories that are not intelligible in isolation but condition and are conditioned by each other in a dialectical whole. The seven categories, compared with their classical formulations, are presented in Table˜4.
@@
Corrêa further identifies social vectors, the specific social forces that carry the potential for transformation, and critiques, the specific structures of domination against which the movement is directed. In my framework, the social vectors are the precariat (the growing class of workers in precarious, platform-mediated employment), the open-source development community (which already practices commons-based peer production at planetary scale), the cooperative movement (from Mondragon to platform cooperatives), the agroecological movement, the community energy movement, and the Fediverse community (which builds and maintains decentralized communication infrastructure). The critiques are directed against techno-feudalism (the concentration of digital productive capacity in a handful of platform corporations), surveillance capitalism (the extraction and monetization of behavioral data), algorithmic management (the replacement of human supervisors with opaque automated systems), fossil-fuel centralization (the structural dependence on concentrated energy sources that enables authoritarian control), and intellectual property enclosure (the artificial scarcification of knowledge, seeds, and code through patent and copyright regimes).
@@
Drawing on Kropotkin’s naturalistic ethics (Kropotkin, 1924), Tolstoy’s principled nonviolence (Tolstoy, 1894), Gandhi’s satyagraha (Gandhi, 1961), and the care ethics tradition (Tronto, 1993), I articulate ten ethical principles that govern the peaceful anarcho-accelerationist project. These principles are not abstract ideals but design constraints for the Liberation Stack, and every technology, protocol, and institution proposed in this paper can be evaluated against them. The first is non-domination: no person, institution, or algorithm shall exercise coercive power over another, and technology must distribute power, never concentrate it. The second is universal care: every system, institution, and technology shall be evaluated by its capacity to sustain life, relationships, and communities, since care is the organizing principle, not profit. The third is nonviolence: social transformation shall be achieved through peaceful means, including technological commons-building, education, mutual aid, prefigurative politics, and democratic persuasion, following Gandhi (Gandhi, 1961) and Sharp (Sharp, 1973). The fourth is commons governance: essential resources such as energy, food, knowledge, communication, code, and productive infrastructure shall be governed as commons according to Ostrom’s principles (Ostrom, 1990), not enclosed as property. The fifth is ecological integrity: all productive activity shall operate within planetary boundaries (Raworth, 2017), since sustainability is a first-order constraint, not an externality. The sixth is radical transparency: algorithms that affect human lives shall be auditable, data collection shall be consensual, and governance shall be open, because opacity is a tool of domination. The seventh is solidarity across difference: following intersectional analysis (Crenshaw, 1989; Collins, 2019), liberation must be universal, since no one is free until everyone is free. The eighth is intellectual commitment: following Chomsky’s principle of the responsibility of intellectuals (Chomsky, 1967), those with knowledge and privilege have a heightened obligation to speak truth to power and to build alternatives. The ninth is freedom of conscience: every person’s spiritual, philosophical, and existential commitments shall be respected, since the framework serves human flourishing in all its dimensions, including the spiritual (Lumbreras and Garrido-Merchán, 2024). The tenth is prefigurative consistency: the means must embody the ends, systems designed for democracy must be governed democratically, and tools for liberation must not reproduce domination (Ward, 1973).
@@
Anarcho-accelerationism rests on four intellectual pillars that bridge the historical and the contemporary. Kropotkin’s Mutual Aid (Kropotkin, 1902) challenged the Social Darwinist misreading of natural selection by documenting extensive cooperation across species. Contemporary evolutionary biology has vindicated this view: multilevel selection theory (Wilson, 2007), reciprocal altruism (Trivers, 1971), and cultural group selection (Henrich, 2016) all demonstrate that cooperation is a primary driver of complex social organization. In computer science, this principle is realized in commons-based peer production (Benkler, 2006). Ostrom’s institutional analysis (Ostrom, 1990) demolished Hardin’s tragedy of the commons by documenting thousands of cases where communities successfully manage shared resources without either privatization or state control. Her eight design principles provide a rigorous institutional framework that maps directly onto distributed system design (De Moor, 2010). Following Stiegler (Stiegler, 2010), I adopt the concept of the pharmakon: every technology is simultaneously remedy and poison. This is not a reason for technological pessimism; it is a reason for technological governance (Feenberg, 2002). Drawing from the anarchist tradition of prefigurative politics (Graeber, 2013), I propose that the systems we build should embody the values they aim to promote.